# Dichography: Two-frame Ultrafast Imaging from a Single Diffraction Pattern

### v2

@@ A.3 Real-space constraint
A similar consideration applies to the retrieval of the support function itself. A fundamental problem in single-particle CDI is the definition of the support function, which is, in most cases, not known a priori. The Shrink-wrap algorithm (SW) [28] has proven to be a fundamental tool for retrieving the correct support function without prior knowledge of the sample. The SW algorithm refines an initially loose support estimate during the reconstruction process, as described in Ref. [28]. Here, the SW algorithm is employed in its original form to independently retrieve the support functions of the two densities, $\rho^{A}$ and $\rho^{B}$. All Dichography results presented in this work used the SW algorithm to determine the samples’ spatial extent, and no a-priori information about the particle shape was employed.
Appendix B Oversampling requirements for Dichography
The oversampling condition is strictly necessary in CDI [32, 6] and it requires the sample density to be surrounded by a region of known scattering density. For single-particle CDI, this condition is enforced by the support function which constraints zero-scattering outside of the sample.
The recorded data gives information only on the amplitude $\left|\psi_{ij}\right|$ of the Fourier representation of the scattering density $\rho$, as the phases are lost. The oversampling condition thus introduces redundancy in the recorded amplitudes which allows them to also encode phase information. For conventional CDI, the amount of information redundancy provided by the recorded intensities $I$ is evaluated through the oversampling degree $O_{d}$, defined as [64]:
$O_{d}$ $=\frac{\text{total number of pixel}}{\text{number of pixel to retrieve}}$ (3)
$=\frac{N\times N}{\sum\textsc{s}_{ij}}$
where $N$ is the linear dimension of the square matrix containing the recorded diffraction data $I$ and s is the binary function that identifies the support. It has values $\textsc{s}_{ij}=0$ for those coordinates where the zero-scattering region is enforced, and $\textsc{s}_{ij}=1$ where the the density values have to be retrieved.
It has been shown that the minimum oversampling requirement for conventional CDI is $O_{d}>2$, to keep the number of “equations” (the number of recorded intensities $N\times N$) higher than the number of “unknowns” (twice the support size, as each retrieved pixel belongs to $\mathbb{C}$). This condition can be interpreted directly in reciprocal space: only the amplitudes of the Fourier representation are accessible as $\sqrt{I_{ij}}$, but twice the amount of information has to be retrieved, i.e. the module $\left|\psi_{ij}\right|$ and the phase $\arg{\left[\psi_{ij}\right]}$.
In the case of Dichography, different $O_{d}$ conditions are required, due to the higher number of unknowns to be retrieved. To better investigate this requirement, it is necessary to re-define the oversampling degree to take into account that two densities $\rho^{A}$ and $\rho^{B}$ are retrieved:
$O^{\prime}_{d}$ $=\frac{\text{total number of pixel}}{\text{number of pixel to retrieve}}$ (4)
$=\frac{N\times N}{\sum\textsc{s}^{A}_{ij}+\sum\textsc{s}^{B}_{ij}}$
where $\textsc{s}^{A}$ and $\textsc{s}^{B}$ are the matrices encoding the support functions of the two frames $\rho^{A}$ and $\rho^{B}$.
As discussed in Sec. A.2, Dichography replaces the squared sum of the two scattering amplitudes with the experimental data $I$. As schematically reported in Fig. 6, for each entry $i,j$ in the matrix, the estimates of the phases for the two fields, $\phi^{A}$ and $\phi^{B}$ are retained. Furthermore, the rescaling operation of the two amplitudes to match the experimental data can be interpreted as a renormalization of a vector, where its phase $\phi^{M}$ is also retained.
Overall, for each recorded intensity $I_{ij}$, three phases have to be retrieved – compared to the single phase value for conventional CDI – $\phi^{A}$ and $\phi^{B}$ and $\phi^{M}$, where the latter encodes the relative amplitude of the two signals. This consideration leads to the conclusion that the minimum condition of the oversampling degree that allows for Dichography is $O_{d}>4$.
By inserting the condition into Eq. (4), the following condition can be derived:
$\sum S^{A}_{ij}+\sum S^{B}_{ij}<\frac{N^{2}}{4}$ (5)
i.e., the sum of the areas of the two supports $\textsc{s}^{A}$ and $\textsc{s}^{B}$ has to be less than a fourth of the whole matrix. For similarly extended frames (i.e. with similar support size), the individual densities $\rho^{A}$ and $\rho^{B}$ cannot extend more than $\nicefrac{{1}}{{8}}$ of the whole matrix area, a significantly more restrictive condition than the $\nicefrac{{1}}{{2}}$ for conventional CDI.
B.1 Uniqueness of the solution
All imaging algorithms for conventional CDI are, in practice, optimization algorithms [32, 7, 40]. They are designed to minimize the difference between the measured experimental data and the Fourier intensities of the reconstructed sample image. Thus, the phase retrieval problem is an optimization problem. The reconstruction $\rho$ produced by phase retrieval algorithms can then be regarded as the solution to the imaging problem. For conventional CDI, it has been demonstrated that the solution is unique – i.e., there exists only one specific sample image that can produce the measured scattering signal – up to a restricted set of ambiguities that can be trivially resolved, such as translation in space or a $180\text{\,}\mathrm{\SIUnitSymbolDegree}$ rotation [7].
The existence of a unique solution for Dichography remains an open question, as a formal mathematical proof is still lacking. We encourage further investigations in this direction, supported by a more rigorous mathematical treatment of the dichographic imaging problem.
Still, we are convinced about the consistency of the imaging results obtained with Dichography. On the one hand, extensive tests on simulated data have never shown cases in which the reconstructed frames exhibited significant deviations from the expected solution – aside from the detrimental effects of noise, as extensively investigated in Sec. S1 of the Supplemental Material [34].
On the other hand, the uniqueness of the solution can be deduced for a specific case – by comparing the dichographic problem with the conventional phase retrieval problem – and it could serve as a starting point for a complete mathematical demonstration. As discussed in Sec. II, the recorded signal $I$ is the incoherent sum of two scattered intensities, i.e., $I=I^{A}+I^{B}$, where $I^{A}$ and $I^{B}$ correspond to the diffraction signals of the densities in the two frames, $\rho^{A}$ and $\rho^{B}$.
If one of the two frames, e.g., $\rho^{A}$, is correctly retrieved, this implies full knowledge of $I^{B}$, as
$I^{B}=I-\left|\mathcal{F}[\rho^{A}]\right|^{2},$ (6)
where $\left|\mathcal{F}[\rho^{A}]\right|^{2}$ is the retrieved intensity $I^{A}$. Eq. (6) reduces to a conventional phase retrieval problem for the frame $\rho^{B}$, whose uniqueness is known. This means that if one frame is correctly retrieved, the other frame has only a unique solution in the dichographic problem.
Appendix C Adaptation of Memetic Phase Retrieval
@@
Single-particle single-shot CDI performed at XFELs comes, however, with additional complications that render the algorithm’s task non-trivial. For example, the central region of the diffraction pattern cannot be recorded due to the presence of a hole in the detector, necessary to prevent damages from the transmitted beam. Furthermore, the brightness of the patterns is most of the time sub-optimal and cannot be increased, as the intensity of the single XFEL pulse is already tuned to its maximum performance. These factors significantly impact the performance of iterative algorithms and render CDI at XFELs considerably more challenging.
@@
Mathematical considerations (see Appendix B) as well as systematic tests on simulated data (see Sec. S1 and Sec. S4 of the Supplemental Material [34]) show that Dichography reconstructions are more demanding than conventional CDI, and that the performance of conventional iterative algorithms is further impacted. Thus, the MPR scheme has been adapted to Dichography, resulting in a significant boost in performance.
The MPR adaptation for Dichography is called Equinox. Equinox has been employed to obtain all imaging results presented in this manuscript, as well as those on simulated data discussed in the Supplemental Material, unless specified otherwise. It is implemented as a Python package directly derived from the SPRING software [40], a recently released toolset for image reconstruction in conventional single-particle single-shot CDI. The open-source implementation of Equinox and its documentation are accessible at https://gitlab.ethz.ch/nux/equinox.
Appendix D Experimental details
D.1 Silver nanoparticles at SwissFEL
The two-color experiment was performed at the Maloja endstation of SwissFEL [68, 69]. Wet-chemically grown silver nanoparticles were injected into the experimental chamber, kept under vacuum conditions, using an electrospray and a set of aerodynamic lenses [70, 71]. The electrospray disperses the liquid containing the dissolved nanoparticles into sub-micron-sized droplets. The nanoparticle concentration is optimized for a single sample per droplet. The droplets flow toward the interaction region along with a buffer gas in a set of aerodynamic lenses, where the liquid evaporates.

### v3

@@ V.1.3 Real-space constraint
A similar consideration applies to the retrieval of the support function itself. A fundamental challenge in single-particle CDI is the definition of the support function, which is, in most cases, not known a priori. The Shrink-wrap algorithm (SW) [28] has proven to be a fundamental tool for retrieving the correct support function without prior knowledge of the sample. The SW algorithm refines an initially loose support estimate during the reconstruction process, as described in Ref. [28]. Here, the SW algorithm is employed in its original form to independently update and retrieve the support functions of the two densities, $\rho^{A}$ and $\rho^{B}$. All Dichography results presented in this work used the SW algorithm to determine the samples’ spatial extent, and no a priori information about the sample geometry (or about the structure of xenon doping for DCDI reconstructions) was employed.
V.2 Further considerations on the dichographic problem
V.2.1 Constraint requirements for Dichography
In this section we mathematically show that, thanks to their structural properties, the samples reconstructed in this work imprint enough information in the overlapping diffraction patterns to allow for their reconstruction.
In particular, a necessary (yet not sufficient) requirement for the existence of a unique solution to the dichographic problem is that the amount of independent information contained in the diffraction data is not smaller than the number of unknowns to retrieve. An effective way to evaluate the ratio between these two quantities is the constraint ratio [35]. Initially introduced in the context of conventional CDI, it has been further extended to the case of the incoherent sum of $N$ diffraction signals [33]. In the specific case of Dichography, i.e., $N=2$, the constraint ratio $\Omega$ can be expressed as:
$\Omega=\frac{1}{2}\frac{\left|\mathcal{A}^{A}\cup\mathcal{A}^{B}\right|}{\left|\textsc{s}^{A}\right|+\left|\textsc{s}^{B}\right|}$ (3)
where $\textsc{s}^{A}$ and $\textsc{s}^{B}$ denote the support functions of the densities $\rho^{A}$ and $\rho^{B}$, while $\mathcal{A}^{A}$ and $\mathcal{A}^{B}$ are the supports of the autocorrelation functions of the densities. The cardinality operation, e.g. $\left|\textsc{s}\right|$, indicates the number of coordinates belonging to the support functions, i.e., their number of pixels in two dimensions. A value of $\Omega>1$ means that the diffraction image contains sufficient information to constrain the density retrieval. Through Eq. (3) it is possible to show that the dichographic problem is at most a factor 2 less constrained than the conventional CDI case [33].
The value of $\Omega$ can be directly calculated via Eq. (3) for the double-hit reconstructions in Sec. II.2, all fulfilling the necessary condition $\Omega>1$. The values for Fig. 4a, Fig. 4b and Fig. 4c are $\Omega=1.22$, $\Omega=1.32$ and $\Omega=1.31$, respectively. While similar, they arise from different geometrical features of the samples [33]. For example, the similarity in size of the support functions in Fig. 4a (which tends to reduce $\Omega$) is compensated for by the distinct overall shapes of the two objects. The $\Omega$ value is similar for Fig. 4c, but due to a different reason. Indeed, here, the two support functions are symmetric, reducing the extension of the individual autocorrelation functions and, thus the information content [33]. A higher value of $\Omega=1.67$ is calculated for Fig. 4d. In fact, the support functions are very different in size and strongly asymmetric for the larger object, increasing the information carried by the autocorrelation. Among the double-hit reconstructions, the one in Fig. 3 is characterized by the smallest value with $\Omega=1.12$. The two support functions are similarly extended and have center-symmetric shapes, rendering the problem less constrained and the reconstruction process more challenging.
The calculation of the constraint ratio for the two-color reconstructions in Fig. 2 is, however, not straightforward. A direct application of Eq. (3), considering only the extension of the autocorrelation and the support of the dopants, provides values of $\Omega=6.49$ and $\Omega=12.65$ for the reconstructions of the patterns in Fig. 2a and Fig. 2d, respectively. While those values are particularly high, due to the sparsity of the xenon agglomerates in the droplet, they represent an underestimate of the actual constraint ratio, as they do not take into account the additional information provided by the interference between the xenon density $\rho_{\chi}$ and the droplet density $\rho_{\bigcirc}$, the latter being known a priori and constrained in the reconstruction. In the following section, we introduce a more rigorous method to calculate the constraint ratio for the DCDI approach [40].
V.2.2 Constraint ratio for Droplet Coherent Diffraction Imaging
For a conventional DCDI reconstruction [40], both the droplet profile $\rho_{\bigcirc}$, known a priori, and the dopant structure $\rho_{\chi}$ contribute to the total scattering density $\rho=\rho_{\bigcirc}+\rho_{\chi}$. Their contribution to the autocorrelation function can thus be separated as:
$\begin{split}A\left[\rho\right]&=\rho\star\rho\\ &=\rho_{\bigcirc}\star\rho_{\bigcirc}+\rho_{\chi}\star\rho_{\chi}+\rho_{\chi}\star\rho_{\bigcirc}+\rho_{\bigcirc}\star\rho_{\chi}\end{split}$ (4)
Among all contributions to the autocorrelation function, the cross-correlation terms $\rho_{\chi}\star\rho_{\bigcirc}$ and $\rho_{\bigcirc}\star\rho_{\chi}$ (i.e., the interference terms between the two densities $\rho_{\bigcirc}$ and $\rho_{\chi}$) and the dopant autocorrelation $\rho_{\chi}\star\rho_{\chi}$ carry information on $\rho_{\chi}$.
The coordinates at which the autocorrelation function effectively provides information on $\rho_{\chi}$ are thus those identified by the autocorrelation support:
$\mathcal{A}_{\bigcirc\chi}=(A\left[\rho\right]-A\left[\rho_{\bigcirc}\right])>0$ (5)
As the information to retrieve corresponds to the support function $\textsc{s}_{\chi}$ of the dopant density $\rho_{\chi}$, the constraint ratio for conventional DCDI follows as:
$\Omega_{\bigcirc\chi}=\frac{\left|\mathcal{A}_{\bigcirc\chi}\right|}{2\,\left|\textsc{s}_{\chi}\right|}$ (6)
It is worth noting that the presence of the droplet increases the information content of the diffraction through the interference terms. Thanks to this holographic information, the imaging problem is significantly more constrained, which readily explains the exceptional convergence properties of the DCDI algorithm observed in previous works [40].
It is now possible to extend the constraint ratio $\Omega_{\bigcirc\chi}$ to the Dichography problem, following the arguments discussed in Ref. [33]. For $N$ scattering contributions, the constraint ratio is evaluated as:
$\Omega_{\bigcirc\chi}=\frac{\left|\bigcup\limits_{n=1}^{N}\mathcal{A}_{\bigcirc\chi}^{n}\right|}{2\,\sum\limits_{n=1}^{N}\left|\textsc{s}_{\chi}^{n}\right|}$ (7)
which reduces for Dichography ($N=2$) to:
$\Omega_{\bigcirc\chi}=\frac{1}{2}\frac{\left|\mathcal{A}_{\bigcirc\chi}^{A}\bigcup\mathcal{A}_{\bigcirc\chi}^{B}\right|}{\left|\textsc{s}_{\chi}^{A}\right|+\left|\textsc{s}_{\chi}^{B}\right|}$ (8)
The constraint ratio adapted to DCDI in Eq. (8), which takes into account the additional information contained in the interference terms, yields values of $\Omega_{\bigcirc\chi}=9.57$ and $\Omega_{\bigcirc\chi}=19.73$ for the reconstructions from Fig. 2a and Fig. 2d, respectively. As expected, those values are higher than those obtained by a direct use of the original formula in Eq. (3).
It is worth noting that the constraint ratio $\Omega$ derives purely from geometrical properties of the samples, and it does not include an evaluation of the quality of the information contained in the diffraction data, in particular concerning the amount of noise due to photon statistics. The inclusion of noise and missing data in the evaluation of the constraint ratio is currently an open problem even in conventional CDI, and it will be investigated in future works. Furthermore, $\Omega>1$ represents a necessary, but in general not sufficient, condition for the uniqueness of the solution.
V.2.3 Non-existence of a conventional CDI solution
In conventional CDI, the failure of a reconstruction attempt is a typical indicator of a double-hit event. This empirical observation raises the question whether this behavior is systematic. Does a conventional CDI solution exist for the incoherent sum of two diffraction patterns? Here, we formally approach this question via polynomial decomposition (following Ref. [70]). We conclude that, in general, such a “fictitious” conventional solution does not exist.
Let $\rho^{A}(\mathbf{x})$ and $\rho^{B}(\mathbf{x})$ be two distinct discrete objects (where $\mathbf{x}$ indicates the discrete representation of the spatial coordinates), with finite support in a $d$-dimensional space ($d\geq 2$). We, thus, exclude the trivial case of identical densities, for which there is no difference between coherent and incoherent sum. The case of an identical physical particle imaged with two different colors is, however, still included, as the same particle is differently sampled, thus corresponding to the case of two different densities.
Their Fourier intensities are given by $I^{A}=\tilde{\rho}^{A}\tilde{\rho}^{A*}$ and $I^{B}=\tilde{\rho}^{B}\tilde{\rho}^{B*}$. We seek a single effective compact object $\rho^{\text{eff}}(\mathbf{x})$ such that:
$\tilde{\rho}^{\text{eff}}(\mathbf{q})\tilde{\rho}^{\text{eff}}(\mathbf{q})^{*}\stackrel{{\scriptstyle?}}{{=}}\tilde{\rho}^{A}(\mathbf{q})\tilde{\rho}^{A}(\mathbf{q})^{*}+\tilde{\rho}^{B}(\mathbf{q})\tilde{\rho}^{B}(\mathbf{q})^{*}\quad\forall\mathbf{q}.$ (9)
The Discrete Fourier Transform (DFT) operation corresponds to the Z-transform evaluated on the unit circle (specifically the unit torus in $d$-dimensions). In particular, for the 2D case, the discrete density function $\rho(\mathbf{x})$ is mapped to a polynomial $P(\mathbf{z})$ via:
$P(z_{1},z_{2})=\sum_{x_{1},x_{2}\in\textsc{s}}\rho(x_{1},x_{2})z_{1}^{-x_{1}}z_{2}^{-x_{2}},$ (10)
where s is the finite support. The diffraction pattern $I$ of an individual density $\rho$ corresponds to the polynomial product $P(\mathbf{z})P^{*}(\nicefrac{{1}}{{\mathbf{z}^{*}}})$ evaluated on the unit circle, i.e., at coordinates $z_{j}=e^{\text{i}k_{j}}$.
Finding a single object $\rho^{\text{eff}}$ requires finding a polynomial $P^{\text{eff}}$ that factorizes the sum of the two diffraction patterns, i.e.:
$P^{\text{eff}}(\mathbf{z}){P^{\text{eff}}}^{*}(\nicefrac{{1}}{{\mathbf{z}^{*}}})=P^{A}(\mathbf{z}){P^{A}}^{*}(\nicefrac{{1}}{{\mathbf{z}^{*}}})+P^{B}(\mathbf{z}){P^{B}}^{*}(\nicefrac{{1}}{{\mathbf{z}^{*}}}).$ (11)
In practical terms, Eq. (11) states that a CDI solution to the Dichography problem, $\rho^{\text{eff}}$, exists only if the sum of the two factorizable polynomials is itself a factorizable polynomial of the form $P^{\text{eff}}(\mathbf{z}){P^{\text{eff}}}^{*}(\nicefrac{{1}}{{\mathbf{z}^{*}}})$.
For the 2D case, and in general for any dimension $d\geq 2$, Hayes proved [70] that the set of factorizable polynomials forms a set of measure zero in the space of all polynomials. Thus, almost always, the sum $P^{A}{P^{A}}^{*}+P^{B}{P^{B}}^{*}$ cannot be expressed as $P^{\text{eff}}{P^{\text{eff}}}^{*}$. This is analogous to the sum of two squared integers, which is almost never a perfect square. As a consequence, the measured intensity $I^{\text{tot}}$ does not correspond to the Fourier transform of any single compact object. The data violates the fundamental consistency constraints of the standard phase retrieval problem (specifically the compactness of the density), and a conventional CDI solution cannot be found. This aspect is further investigated numerically in Sec. 6 of the Supplemental Material, along with the discussion of pathological cases for which, due to strict symmetry features, a conventional CDI solution indeed exists.
V.2.4 Uniqueness of the solution
A fundamental aspect of optimization problems is the uniqueness of the optimal solution. Are there equivalent solutions to the dichographic optimization problem? Specifically, are the two densities $\rho^{A}$ and $\rho^{B}$ retrieved by Dichography unique? While Hayes’ theorem [70] guarantees the uniqueness of phase retrieval for a single compact object in dimensions $d\geq 2$, the incoherent sum of two intensities introduces an additional degree of freedom. And, indeed, it is easy to verify that a rotation (or, more generally, a unitary transform [33]) of the densities produces the same diffraction pattern. Specifically, it is possible to create a new pair of frames $\rho^{\prime A}$ and $\rho^{\prime B}$ via the linear operation:
$\begin{pmatrix}\rho^{\prime A}(\mathbf{x})\\ \rho^{\prime B}(\mathbf{x})\end{pmatrix}=\begin{pmatrix}\cos\theta&\sin\theta\\ -\sin\theta&\cos\theta\end{pmatrix}\begin{pmatrix}\rho^{A}(\mathbf{x})\\ \rho^{B}(\mathbf{x})\end{pmatrix}$. (12)
These two densities fully satisfy the experimental constraint $|\tilde{\rho}^{\prime A}|^{2}+|\tilde{\rho}^{\prime B}|^{2}=I^{\text{tot}}$. Here, the range of angles for which the ambiguity is relevant is restricted to $\theta\in\left[0,\pi/2\right)$. The other regions of possible rotations can be easily related to this interval by combining the operation of frame swapping, under which Dichography is inherently ambiguous, or the assignment of a global sign to the density values, under which even conventional phase retrieval is ambiguous.
Mathematically, the diffraction data alone cannot distinguish the true densities from a rotation in Eq. (12). There are, however, physical constraints that break this rotational equivalence, rendering the solution unique (up to the ambiguities that are also affecting conventional phase retrieval). A key aspect is the role of the support functions in combination with the constraint of a positive density, known as positivity constraint [34]. Given $\textsc{s}^{A}$ and $\textsc{s}^{B}$, we can identify three possible cases:
Incompatible supports
In this case $\textsc{s}^{A}\neq\textsc{s}^{B}$ and neither support contains the other. Only the rotation $\theta=0$ renders the reconstructions compatible with their own support function. Any other rotation would require a support function $\textsc{s}=\textsc{s}^{A}\cup\textsc{s}^{B}$ for both densities. This is further enhanced by the positivity constraint. Indeed, any rotation apart from $\theta=0$ would produce a reconstruction with negative density values, forbidden by the constraint.
Nested supports
In this case, $\textsc{s}^{A}\neq\textsc{s}^{B}$ but $\textsc{s}^{A}\subset\textsc{s}^{B}$. Again, the rotational symmetry is broken, but in a “milder” way. Indeed, while a rotation would render $\rho^{\prime A}$ incompatible with $\textsc{s}^{A}$, $\rho^{\prime B}$ would still be compatible with $\textsc{s}^{B}$, as $\textsc{s}^{B}=\textsc{s}^{A}\cup\,\textsc{s}^{B}$. In this case, the positivity constraint doesn’t necessarily help as there can be rotations that do not yield negative values. The same considerations apply by swapping the frames $A$ and $B$.
Identical supports
In this case $\textsc{s}^{A}=\textsc{s}^{B}$ and the rotational ambiguity cannot be resolved by the support functions. A unique solution can be identified thanks to the sole positivity constraint under specific conditions [33].
It is worth noting that, even in the case of two-color pulses hitting a single particle that retains the shape during the two diffraction events, the support functions of the two reconstructed frames have different spatial extensions. Indeed, the two pulses sample the particle at different resolutions, producing two different discrete representations of the reconstructed densities (see Sec. V.6). Thus, the incompatibility of the support functions that resolves the rotational ambiguity is inherently provided by two-color pulses (i.e., two-color data generally cannot fall into the tricky “identical supports” case).
The discussion so far has been based on the a priori knowledge of the support functions $\textsc{s}^{A}$ and $\textsc{s}^{B}$. In typical cases, those support functions are not known and they are retrieved during the reconstruction process with the Shrink-wrap algorithm [28]. In theory, the Shrink-wrap algorithm doesn’t prevent the possibility to reach two support functions $\textsc{s}^{\prime A}=\textsc{s}^{\prime B}=\textsc{s}^{A}\cup\,\textsc{s}^{B}$, thus restoring the rotational ambiguity. However, in practical cases, we observe that the Shrink-wrap tends to minimize the extension of the two support functions, i.e., tends to achieve the lowest possible spatial extension of the two frames. This action of the Shrink-wrap naturally pushes the reconstruction process away from the $\textsc{s}^{A}\cup\textsc{s}^{B}$ case, thus resolving the ambiguity in both the “Incompatible supports” and “Nested supports” cases.
The rotational ambiguity is, thus, theoretically resolved by the support constraint. For “ideal” data, not affected by noise, the only zero-error solution is the one where $\theta=0$ and the two frames are perfectly separated. Real experimental diffraction patterns are, however, affected by shot noise, which renders even the correct solution not fully compatible with the constraints. This leaves “room” for other reconstructions close to the real solution to arise, similarly to what happens with the twin-image problem in conventional CDI [71]. In the case of Dichography, those reconstructions close-by in the error landscape are represented by slight rotations in Eq. (12). We believe that these small, unresolved rotations are the origin of the ghosting artifacts observed in the experimental reconstructions, where a faint “ghost” of one frame appears in the other.
V.3 Adaptation of Memetic Phase Retrieval
@@
Single-particle single-shot CDI performed at XFELs, however, comes with additional complications that render the algorithm’s task non-trivial. For example, the central region of the diffraction pattern cannot be recorded due to the presence of a hole in the detector, necessary to prevent damage from the transmitted beam. Furthermore, the brightness of the patterns is most of the time sub-optimal and cannot be increased, as the intensity of the single XFEL pulse is already tuned to its maximum performance. These factors significantly impact the performance of iterative algorithms and render CDI at XFELs considerably more challenging.
@@
Mathematical considerations (see Sec. V.2.1) as well as systematic tests on simulated data (see Sec. S1 and Sec. S4 of the Supplemental Material [39]) show that Dichography reconstructions are more demanding than conventional CDI, and that the performance of conventional iterative algorithms is further impacted. Thus, the MPR scheme has been adapted to Dichography resulting in a significant boost in performance, especially concerning the retrieval of the support function during the reconstruction process.
The MPR adaptation for Dichography is called Equinox. Equinox has been employed to obtain all imaging results presented in this manuscript, as well as those on simulated data discussed in the Supplemental Material, unless specified otherwise. It is implemented as a Python package directly derived from SPRING [45], a recently released toolset for image reconstruction in conventional single-particle single-shot CDI. The open-source implementation of Equinox and its documentation are accessible at https://gitlab.ethz.ch/nux/equinox.
Equinox retains the conceptual structure of its conventional CDI counterpart, SPRING [45]. The most significant modification, aside from the use of dichographic iterative projectors (see Sec. V.1), lies in the initialization of the starting densities. As in SPRING, the individual frames are initialized by populating the real-space matrix with spherical density profiles of random position and size. In Equinox, however, the densities of the two frames are additionally renormalized to reflect unequal scattering contributions. The intensity ratio is randomly assigned within a range from 50:50 to 90:10. While these limits are arbitrary and tunable, we empirically observed that this range is sufficiently robust to reconstruct data with diverse intensity ratios without the need for case-specific fine-tuning. The algorithm can successfully retrieve the correct scattering ratio independent of the initial balancing.
Currently, Equinox implements the Error Reduction (ER), Hybrid Input-Output (HIO) [36], and Relaxed Averaged Alternating Reflections (RAAR) [75] algorithms. Future work includes the implementation of the Difference Map (DM) algorithm [37] adapted for the dichographic problem [33], which is expected to further improve the convergence properties of the solver.
V.4 Experimental details
V.4.1 Silver nanoparticles at SwissFEL
The experiment was performed at the Maloja endstation of SwissFEL [76, 77]. Wet-chemically grown silver nanoparticles were injected into the experimental chamber, kept under vacuum conditions, using an electrospray and a set of aerodynamic lenses [78, 79]. The electrospray disperses the liquid containing the suspended nanoparticles into sub-micron-sized droplets. The nanoparticle concentration is optimized for a single sample per droplet. The droplets flow toward the interaction region along with a buffer gas in a set of aerodynamic lenses, where the liquid evaporates.